# Revealing isotropic abundant low-energy excitations in UTe2 through complex microwave surface impedance

### v1

@@ I.1 Experiment
We perform cavity perturbation experiments with these samples in a cylindrical, hollow rutile dielectric resonator. A microwave signal is sent through the cavity, and the transmission is measured, which is shown in Fig. 1. The sample perturbs the resonances of the cavity revealing the electrodynamic properties through changes in the resonant frequency $f_{0}$ and quality factor $Q$ of the modes as functions of sample temperature. Resonator perturbation experiments like this have been used to study the electrodynamics of a diverse variety of superconductors for many years [21, 22, 23, 24, 25, 26, 27, 28, 7]. A novel aspect of this work is that numerous modes in the spectrum shown in Fig. 1 are utilized to make multiple independent measurements of the sample surface impedance. Another novelty of this work is the use of the anisotropic normal state resistivity tensor to determine the two parameters ($G$ and $X_{0}$) that convert the raw data into the absolute surface impedance and magnetic penetration depth.
@@ I.2 Resonator
The sample is introduced into a hollow cylindrical, rutile dielectric resonator [7] to perturb its resonances. The sample is attached to the end of a sapphire-rod ‘hot finger’ [22, 23, 26, 27] using Apiezon N-grease, which holds it on the symmetry axis in the middle of the resonator and allows heat to be conducted from a heater outside the resonator cavity directly to the sample over a range from 100 mK to 20 K, while isolating it from the copper walls of the cavity and the rutile, which are nominally held at 100 mK. The microwave transmission through the resonator is measured for frequencies between 2 and 12 GHz, and the resonant frequency $f_{0}$ and quality factor $Q$ of many modes are determined by fitting [29]. The resonant properties are measured both with and without the sample present in the cavity, which allows us to isolate an effective $\Delta f_{0}$ and $Q$ arising from the sample alone [26, 7], where $\Delta f_{0}$ denotes the change in $f_{0}$ from its value at the lowest measurement temperature $T_{0}\sim 150$ mK. See §A.2 for futher details.
@@ II Data analysis (utilizing minimal assumptions)
The first electrodynamic property of interest is the complex surface impedance $Z_{s}=R_{s}+iX_{s}$, made up of the surface resistance $R_{s}(T)$ and surface reactance $X_{s}(T)$. Note that the surface impedance which we measure is a composite of multiple anisotropic responses. Our model of this is shown in Eq. 4, and will be discussed in detail in §III.2 below. We recover the real and imaginary components of the surface impedance from the sample resonance data as functions of temperature [26, 7],
$Z_{s}=\frac{G}{Q}+i\left[-2G\frac{\Delta f_{0}}{f_{0,\text{tot.}}(T_{0})}+X_{0}\right].$ (1)
Here $X_{0}$ is the minimum-temperature surface reactance, which sets the starting point for $X_{s}$ since this measurement is only directly sensitive to relative changes in $X_{s}$. The geometry factor $G$ (Eq. 9) has dimensions of resistance and relates the field structure at the sample location to that of the rest of the resonator for each mode. Here we assume that the sample creates a predominantly magnetic perturbation to the resonant modes considered. See §A.2 for a discussion of the perturbation due to the sample. To fully relate the measured $f_{0}$ and $Q$ to $Z_{s}$, we must determine $G$ and $X_{0}$, which are unique to each mode of the resonator, as well as the sample and its orientation.
@@
Figure 2: Plots of change in penetration depth $\Delta\lambda/G=(\lambda-\lambda_{0})/G$ vs. temperature and power-law fits for $T\leq T_{c}/3$. The different colors correspond to different values of $X_{0}/G$ over its allowed range, from minimum (red) to maximum (purple). The insets show the underlying composite dimension-less surface impedance $Z_{s}(T)/G$ data for the same $X_{0}/G$ values as used in the $\Delta\lambda$ plots. The black curve in the insets is $R_{s}(T)/G$ as dictated by the $Q(T)$ data while the colored curves are possible $X_{s}(T)/G$ as dictated by the $f_{0}(T)$ data and $X_{0}/G$ choice.
In the local limit, the surface impedance is related to the complex conductivity $\sigma=\sigma_{1}-i\sigma_{2}$ by $Z_{s}=\sqrt{\frac{i\mu_{0}\omega}{\sigma}}$, where $\omega=2\pi f$. We discuss our results for the complex conductivity in §III.3 below. When $\sigma_{2}\gg\sigma_{1}$, the penetration depth is given by, $\lambda=1/\sqrt{\mu_{0}\omega\sigma_{2}}$ [7, 32]. Fig. 2 shows the temperature dependence well below the microwave critical temperature, $T_{c}$, of the change in the penetration depth from its zero temperature value $\lambda_{0}$, for two UTe2 samples, using only the raw $f_{0}(T)$ and $Q(T)$ data for multiple possible values of $X_{0}/G$. Importantly, the $\Delta\lambda(T)/G$ data demonstrates an independence on $X_{0}/G$ in this temperature range.
We also perform power-law fits to this penetration depth data for $T\leq T_{c}/3$ which are included in Fig. 2 as dashed lines. See §D.1.1 for a discussion of this fitting. A power-law temperature dependence for the penetration depth at low temperatures can be indicative of nodal superconductivity. The form,
@@
has theoretical predictions for the dimensionless parameters $\alpha$ and $\eta$ depending on the pairing symmetry, direction of current flow, and scattering conditions present in the sample [33, 34]. For both UTe2 B39 and UTe2 B40, we estimate $T_{c}=1.96$ K. We used data from base temperature to $T_{c}/3$ to do the power law fitting in Eq. 2 with $\lambda_{0}/G$, $\alpha$, and $\eta$ as the fitting parameters for the composite $\lambda/G$ data. We also compare this low-temperature $\lambda(T)$ for UTe2 with that of s-wave superconductors Nb and NbSe2 measured in the same apparatus in §D.1.3.
@@
for $T\leq T_{c}/3$, which we use to determine the zero-temperature surface resistance $R_{0}$ for each mode. In fact, the determination of $\alpha_{R}$, $\eta_{R}$, and $R_{0}/G$ is explicitly independent of $G$ and $X_{0}$. See §D.2 for a discussion of this fitting. Similarly to the penetration depth power-law exponents, the surface resistance power-law exponents, $\alpha_{R}$, show no systematic variation between modes. These give average results of $\alpha_{R}=1.70\pm 0.20$ for UTe2 B39 and $\alpha_{R}=1.86\pm 0.20$ for UTe2 B40, which are effectively the same as the penetration depth power-law exponents considering the uncertainties. Table LABEL:tab:lambda_and_loss also includes a summary of $\alpha_{R}$ and the nominally frequency-independent residual loss $R_{0}/\omega^{2}$. The significance of this power-law behavior is not straightforward to interpret; see §D.2 for more discussion. We note in passing that a power-law behavior of $R_{s}(T)$ with $\alpha_{R}\approx 1$ independent of frequency was observed in CeCoIn5 [28].
@@ III.2 Impedance fitting
where the weights $w_{i}$ satisfy $0\leq w_{i}\leq 1$ and $\Sigma w_{i}=1$ [38]. These weights represent the projections of the surface currents along the crystallographic axes. By studying a number of diverse modes with different projection sets, a sampling of $w_{i}$ is obtained, giving information about the full surface impedance tensor.
@@
We use the measured anisotropic DC resistivity tensor of similar UTe2 crystals measured with a Montgomery geometry [39] to do this fitting. See §B.1 for further discussion of our treatment of this resistivity data. The normal-state surface impedance fitting is performed from 2 to 20 K for each mode, individually, which determines the fitting parameters $G$, $X_{0}$, and the $w_{i}$ for each mode. The results of this fitting for a representative mode are shown in Fig. 5. This fitting and its uncertainties are discussed in detail in §B.3. For this mode, we obtain fitting parameters $G=11.2\pm 0.5\ k\Omega$, $X_{0}=96\pm 13\ m\Omega$, $w_{a}=0.35\pm 0.12$, $w_{b}=0.17\pm 0.10$, and $w_{c}=0.49\pm 0.02$. Note that the uncertainty of $w_{b}$ is comparable to the value of $w_{b}$ itself, indicating that there may not be a significant response of currents corresponding to the b-axis for this mode.
@@ IV Discussion/ Interpretation
The results for $\alpha$ and $\eta$ shown in Figs. 9(a), 9(b) are inconsistent with a single pair of point nodes (simple p-wave) in the weak scattering limit. We would expect the data to be somewhere between the parallel ($\alpha=2$) and perpendicular ($\alpha=4$) predictions with only a few modes close to the direction of the nodes matching the parallel-to-nodes results, while many more modes would produce the perpendicular-to-nodes results since the directions perpendicular to the nodes form a line rather than a point in phase space. These results are also inconsistent with a single line node in the weak scattering limit, for which $\alpha=1$ and $\alpha=3$ are expected parallel and perpendicular to the line node, respectively [34]. However, our data could be consistent with either a line node or a pair of point nodes in the case of unitary scattering. For both these cases, $\alpha\approx 2$ is expected. Further, $\eta\approx 0.5-0.7$ is expected for point nodes with unitary scattering, and $\eta\approx 0.8$ is expected for a line node with unitary scattering, [34], both of which are consistent with our data. The presence of multiple pairs of point nodes slightly off of one or more of the high symmetry axes could also be consistent with these results, which is used to explain the results of [8]. We note that a nearly isotropic $\Delta\lambda\sim T^{1.9-2}$ was observed in PrOs4Sb12 and attributed to multi-domain orientations of a point nodal order parameter [42].
@@
Another possible explanation for the behavior seen in Fig. 9(a) is the presence of Weyl nodes in a topological superconducting state. Although broken time-reversal invariance is not observed in recent generation UTe2 samples, theory suggests that different irreducible representations (e.g. B2u and B3u) may be nearly degenerate, leading to a mixed order parameter [45]. Alternatively, recent theory suggests that a combined Kondo/spin-liquid model can show a single-step-transition into a chiral topological superconducting state [46]. The theory for a topological Weyl superconductor predicts the existence of a gapless 2D Majorana fermion surface fluid with unique electrodynamic properties. First, a Weyl node superconductor is expected to show power-law exponent $\alpha=2$ for the penetration depth for currents in all directions [47], which is qualitatively consistent with our data and that of [7, 8], though some measured values of $\alpha$ are considerably less than 2. Secondly, a new absorption mechanism is predicted to exist involving absorption of a photon from a surface state causing a transition to an empty bulk state. This mechanism is predicted to be active at zero temperature and to show a characteristic anisotropy, being larger for electric fields polarized parallel to the Weyl nodes compared to the perpendicular orientation. This result is predicted to be robust to disorder as well [47].
@@ VI Acknowledgments
The work of AC-J and SMA was supported by NSF-DMR/2004386, and ARO/FSDL under grant W911NF-24-1-0153. The work of AS, YSE, IMH, SRS, JP and NPB was supported by by the U.S. Department of Energy Award No. DE-SC-0019154 (sample characterization), the Air Force Office of Scientific Research under Grant No. FA9950-22-1-0023 (materials synthesis), and the Gordon and Betty Moore Foundation’s EPiQS Initiative through Grant No. GBMF9071, the National Institute of Standards and Technology, and the Maryland Quantum Materials Center. SRS acknowledges support from the National Institute of Standards and Technology Cooperative Agreement 70NANB17H301.
@@ A.2 Selection of perturbed resonant cavity modes
$\begin{split}\Delta f_{0}(T)&=\Delta f_{0,\text{tot.}}(T)-\Delta f_{0,r}(T)\\ 1/Q(T)&=1/Q_{\text{tot.}}(T)-1/Q_{r}(T),\end{split}$ (8)
@@ B.2 Surface Reactance constraints
Figure 13: Complex geometry of the (a) surface impedance $Z_{s}=R_{s}+iX_{s}$ and (b) complex conductivity $\sigma=\sigma_{1}-i\sigma_{2}$, where $\lambda_{L}$ is the London penetration depth and $\omega$ is angular frequency. The axes are dimensionless versions of (a) $R_{s},X_{s}$, and (b) $\sigma_{1},\sigma_{2}$. Curves of constant superfluid fraction $f_{s}=0,0.2,0.4,0.6,\text{and}\ 0.8$ are shown in red and constant scattering time $\omega\tau=0.2,0.4,0.6,0.8,\text{and}\ 1.0$ are shown in blue. The lower dashed line in (a) shows the condition $R_{s}=X_{s}$, and the upper dashed line shows $R_{s}=\sqrt{3}X_{s}$. To find the direction of increasing $f_{s}$, follow any curve of constant $\omega\tau$ towards $(0,1)$ and vice versa.

### v2

@@ I.1 Experiment
We perform cavity perturbation experiments with these samples in a cylindrical, hollow rutile dielectric resonator. A microwave signal is sent through the cavity, and the transmission is measured, which is shown in Fig. 1. The sample perturbs the resonances of the cavity revealing the electrodynamic properties through changes in the resonant frequency $f_{0}$ and quality factor $Q$ of the modes as functions of sample temperature. Resonator perturbation experiments like this have been used to study the electrodynamics of a diverse variety of superconductors for many years [21, 22, 23, 24, 25, 26, 27, 28, 7]. A novel aspect of this work is that numerous modes in the spectrum shown in Fig. 1 are utilized to make multiple independent measurements of the sample surface impedance. Another novelty of this work is the use of the anisotropic normal state resistivity tensor to determine the two parameters ($G$ and $X_{0}$) that convert the raw data into the absolute surface impedance and magnetic penetration depth. Both of these novel features are discussed in detail in §A and §B.
@@ I.2 Resonator
The sample is introduced into a hollow cylindrical, rutile dielectric resonator [7] to perturb its resonances. The sample is attached to the end of a sapphire-rod ‘hot finger’ [22, 23, 26, 27] using Apiezon N-grease, which holds it on the symmetry axis in the middle of the resonator and allows heat to be conducted from a heater outside the resonator cavity directly to the sample over a range from 100 mK to 20 K, while isolating it from the copper walls of the cavity and the rutile, which are nominally held at 100 mK. The microwave transmission through the resonator is measured for frequencies between 2 and 12 GHz, and the resonant frequency $f_{0}$ and quality factor $Q$ of many modes are determined by fitting [29]. The resonant properties are measured both with $(f_{0,\text{tot.}},Q_{\text{tot.}})$ and without $(f_{0,r},Q_{r})$ the sample present in the cavity, which allows us to isolate an effective $\Delta f_{0{\color[rgb]{0,0,0},\text{sample}}}$ and $Q_{{\color[rgb]{0,0,0}\text{sample}}}$ arising from the sample alone [26, 7], where $\Delta f_{0}$ denotes the change in $f_{0}$ from its value at the lowest measurement temperature $T_{0}\sim 150$ mK. See §A.2 for further details.
@@ II Data analysis (utilizing minimal assumptions)
The first electrodynamic property of interest is the complex surface impedance $Z_{s}=R_{s}+iX_{s}$, made up of the surface resistance $R_{s}(T)$ and surface reactance $X_{s}(T)$. Note that the surface impedance which we measure is a composite of multiple anisotropic responses , namely the 3 components of the diagonal surface impedance tensor. Our model of this is shown in Eq. 4, and will be discussed in detail in §III.2 below. We recover the real and imaginary components of the surface impedance from the sample resonance data as functions of temperature [26, 7],
$Z_{s}=\frac{G}{Q_{{\color[rgb]{0,0,0}\text{sample}}}}+i\left[-2G\frac{\Delta f_{0{\color[rgb]{0,0,0},\text{sample}}}}{f_{0,\text{tot.}}(T_{0})}+X_{0}\right].$ (1)
Here $X_{0}$ is the minimum-temperature surface reactance, which sets the starting point for $X_{s}$ since this measurement is only directly sensitive to relative changes in $X_{s}$. The geometry factor $G$ (Eq. 9) has dimensions of resistance and relates the field structure at the sample location to that of the rest of the resonator for each mode. Here we assume that the sample creates a predominantly magnetic perturbation to the resonant modes considered. See §A.2 for a discussion of the perturbation due to the sample. To fully relate the measured $f_{0{\color[rgb]{0,0,0},\text{sample}}}$ and $Q_{{\color[rgb]{0,0,0}\text{sample}}}$ to $Z_{s}$, we must determine $G$ and $X_{0}$, which are unique to each mode of the resonator, as well as the sample and its orientation.
@@
Figure 2: Plots of reduced penetration depth $\lambda/G$ vs. an intermediate exponent of temperature from power-law fits (shown as dashed lines) for $T\leq T_{c}/3$ for (a) sample B39 and (b) sample B40. The different colors correspond to different values of $X_{0}/G$ over its allowed range, from minimum (red) to maximum (purple), for which the power-law $\alpha$ slightly changes. The insets show the change in penetration depth $\Delta\lambda/G=(\lambda-\lambda_{0})/G$ with the same respective x-axes. These plots largely overlap for each value of $X_{0}/G$.
In the local limit, the surface impedance is related to the complex conductivity $\sigma=\sigma_{1}-i\sigma_{2}$ by $Z_{s}=\sqrt{\frac{i\mu_{0}\omega}{\sigma}}$, where $\omega=2\pi f$. We discuss our results for the complex conductivity in §III.3 below. When $\sigma_{2}\gg\sigma_{1}$, the penetration depth is given by, $\lambda=1/\sqrt{\mu_{0}\omega\sigma_{2}}$ [7, 32]. Both $\sigma$ and $\lambda$ calculated from composite $Z_{s}$ in this way are also composite quantities. Fig. 2 shows the temperature dependence well below the microwave critical temperature, $T_{c}$, of the penetration depth, for two UTe2 samples, using only the raw ${\color[rgb]{0,0,0}\Delta}f_{0{\color[rgb]{0,0,0},\text{sample}}}(T)$ and $Q_{{\color[rgb]{0,0,0}\text{sample}}}(T)$ data for multiple possible values of $X_{0}/G$. Importantly, the $\Delta\lambda(T)/G$ data , shown in the insets of Fig. 2, demonstrates an independence on $X_{0}/G$ in this temperature range.
A power-law temperature dependence for the penetration depth at low temperatures ($T\leq T_{c}/3$) can be indicative of nodal superconductivity. The form,
@@
has theoretical predictions for the dimensionless parameters $\alpha$ and $\eta$ depending on the pairing symmetry, direction of current flow, and scattering conditions present in the sample [33, 34]. For both UTe2 B39 and UTe2 B40, we estimate $T_{c}=1.96$ K. We used data from base temperature to $T_{c}/3$ to do the power law fitting in Eq. 2 with $\lambda_{0}/G$, $\alpha$, and $\eta$ as the fitting parameters for the composite $\lambda/G$ data. These power-law fits are included in Fig. 2 as dashed lines. See §D.1.1 for a detailed discussion of this fitting, as well as Table 3 which summarizes the uncertainties of the parameters in this fit and others. We also compare this low-temperature $\lambda(T)$ for UTe2 with that of s-wave superconductors Nb and NbSe2 measured in the same apparatus in §D.1.3.
@@
for $T\leq T_{c}/3$, which we use to determine the zero-temperature surface resistance $R_{0}$ for each mode. In fact, the determination of $\alpha_{R}$, $\eta_{R}$, and $R_{0}/G$ is explicitly independent of $G$ and $X_{0}$. See §D.2 for a detailed discussion of this fitting , as well as Table 3 which summarizes the uncertainties of the parameters in this fit and others. Similarly to the penetration depth power-law exponents, the surface resistance power-law exponents, $\alpha_{R}$, show no systematic variation between modes. These give average results of $\alpha_{R}=1.70\pm 0.20$ for UTe2 B39 and $\alpha_{R}=1.86\pm 0.20$ for UTe2 B40, which are effectively the same as the penetration depth power-law exponents considering the uncertainties. Table LABEL:tab:lambda_and_loss also includes a summary of $\alpha_{R}$ and the nominally frequency-independent residual loss $R_{0}/\omega^{2}$. The significance of this power-law behavior is not straightforward to interpret; see §D.2 for more discussion. We note in passing that a power-law behavior of $R_{s}(T)$ with $\alpha_{R}\approx 1$ independent of frequency was observed in CeCoIn5 [28].
@@ III.2 Impedance fitting
where the weights $w_{i}$ satisfy $0\leq w_{i}\leq 1$ and $\Sigma w_{i}=1$ [38]. These weights represent the projections of the surface currents along the crystallographic axes. By studying a number of diverse modes with different projection sets, a sampling of $w_{i}$ is obtained, giving information about the full surface impedance tensor. Other composite quantities, such as complex conductivity $\sigma$ and penetration depth $\lambda$ would in general have a more complicated dependence on their tensor components. Certain simplifying limits are still possible; such as, at low temperatures ($T\ll T_{c}$), $\lambda$ would decompose linearly with the same current induced weights $w_{i}$ as in Eq. 4 since $X_{s}\approx\mu_{0}\omega\lambda$ in this limit as discussed in §D.4. This is demonstrated for the zero-temperature penetration depth $\lambda_{0}$ in Fig. 21(b).
@@
We use the measured anisotropic DC resistivity tensor of similar UTe2 crystals measured with a Montgomery geometry [39] to do this fitting. See §B.1 for further discussion of our treatment of this resistivity data. The normal-state surface impedance fitting is performed from 2 to 20 K for each mode, individually, which determines the fitting parameters $G$, $X_{0}$, and the $w_{i}$ for each mode. The results of this fitting for a representative mode are shown in Fig. 5. This fitting and its uncertainties are discussed in detail in §B.3 , and summarized in Table 3. For this mode, we obtain fitting parameters $G=11.2\pm 0.5\ k\Omega$, $X_{0}=96\pm 13\ m\Omega$, $w_{a}=0.35\pm 0.12$, $w_{b}=0.17\pm 0.10$, and $w_{c}=0.49\pm 0.02$. Note that the uncertainty of $w_{b}$ is comparable to the value of $w_{b}$ itself, indicating that there may not be a significant response of currents corresponding to the b-axis for this mode.
@@ IV Discussion/ Interpretation
The results for $\alpha$ and $\eta$ shown in Figs. 9(a), 9(b) are inconsistent with a single pair of point nodes (simple p-wave) in the weak scattering limit. We would expect the data to be somewhere between the parallel ($\alpha=2$) and perpendicular ($\alpha=4$) predictions with only a few modes close to the direction of the nodes matching the parallel-to-nodes results, while many more modes would produce the perpendicular-to-nodes results since the directions perpendicular to the nodes form a line rather than a point in phase space. These results are also inconsistent with a single line node in the weak scattering limit, for which $\alpha=1$ and $\alpha=3$ are expected parallel and perpendicular to the line node, respectively [34]. However, our data could be consistent with either a line node or a pair of point nodes in the case of unitary scattering. For both these cases, $\alpha\approx 2$ is expected. Further, $\eta\approx 0.5-0.7$ is expected for point nodes with unitary scattering, and $\eta\approx 0.8$ is expected for a line node with unitary scattering, [34], both of which are consistent with our data. The presence of multiple pairs of point nodes slightly off of one or more of the high symmetry axes could also be consistent with these results [42], which is used to explain the results of [8]. We note that a nearly isotropic $\Delta\lambda\sim T^{1.9-2}$ was observed in PrOs4Sb12 and attributed to multi-domain orientations of a point nodal order parameter [43].
@@
Another possible explanation for the behavior seen in Fig. 9(a) is the presence of Weyl nodes in a topological superconducting state. Although broken time-reversal invariance is not observed in recent generation UTe2 samples, theory suggests that different irreducible representations (e.g. B2u and B3u) may be nearly degenerate, leading to a mixed order parameter [46]. Alternatively, recent theory suggests that a combined Kondo/spin-liquid model can show a single-step-transition into a chiral topological superconducting state [47]. The theory for a topological Weyl superconductor predicts the existence of a gapless 2D Majorana fermion surface fluid with unique electrodynamic properties. First, a Weyl node superconductor is expected to show power-law exponent $\alpha=2$ for the penetration depth for currents in all directions [48], which is qualitatively consistent with our data and that of [7, 8], though some measured values of $\alpha$ are considerably less than 2. Secondly, a new absorption mechanism is predicted to exist involving absorption of a photon from a surface state causing a transition to an empty bulk state. This mechanism is predicted to be active at zero temperature and to show a characteristic anisotropy, being larger for electric fields polarized parallel to the Weyl nodes compared to the perpendicular orientation. This result is predicted to be robust to disorder as well [48]. Predictions of surface states in UTe2 due to Dirac Majorana modes have also been put forward [49, 42]. Finally, recent STM measurements on UTe2 show evidence of zero-energy Andreev conductance on the $(0\bar{1}1)$ surface termination [50].
@@ VI Acknowledgments
The work of AC-J and SMA was supported by NSF-DMR/2004386, and ARO/FSDL under grant W911NF-24-1-0153. The work of AS, YSE, IMH, SRS, JP and NPB was supported by the U.S. Department of Energy Award No. DE-SC-0019154 (sample characterization), the Air Force Office of Scientific Research under Grant No. FA9950-22-1-0023 (materials synthesis), and the Gordon and Betty Moore Foundation’s EPiQS Initiative through Grant No. GBMF9071, the National Institute of Standards and Technology, and the Maryland Quantum Materials Center. SRS acknowledges support from the National Institute of Standards and Technology Cooperative Agreement 70NANB17H301.
@@ A.2 Selection of perturbed resonant cavity modes
$\begin{split}\Delta f_{0{\color[rgb]{0,0,0},\text{sample}}}(T)&=\Delta f_{0,\text{tot.}}(T)-\Delta f_{0,r}(T)\\ 1/Q_{{\color[rgb]{0,0,0}\text{sample}}}(T)&=1/Q_{\text{tot.}}(T)-1/Q_{r}(T),\end{split}$ (8)
@@ B.2 Surface Reactance constraints
Figure 13: Complex geometry of the (a) surface impedance $Z_{s}=R_{s}+iX_{s}$ and (b) complex conductivity $\sigma=\sigma_{1}-i\sigma_{2}$, where $\lambda_{L}$ is the London penetration depth and $\omega$ is angular frequency. The axes are dimensionless versions of (a) $R_{s},X_{s}$, and (b) $\sigma_{1},\sigma_{2}$. Curves of constant superfluid fraction $f_{s}=0,0.2,0.4,0.6,\text{and}\ 0.8$ are shown in red and constant scattering time $\omega\tau=0.2,0.4,0.6,0.8,\text{and}\ 1.0$ are shown in blue. The lower dashed line in (a) shows the condition $X_{s}=R_{s}$, and the upper dashed line shows $X_{s}=\sqrt{3}R_{s}$. To find the direction of increasing $f_{s}$, follow any curve of constant $\omega\tau$ towards $(0,1)$ and vice versa.